OpenStudy (aryavegapunk):

The height of a thrown object as a function of the horizontal distance that it has traveled is a quadratic function. An object is thrown from ground level, 63 m away from the base of a tower 17 m high and just barely clears the top of the tower. it lands 63 m on the other side. Find the quadratic function that represents this. My work so far: h(d)=ad2 + bd + c 17=(63)2a + (63)b 0=(126)2a + 126b 17=(3969)a + 63b 0= 15876a + 126b after this i'm stuck

OpenStudy (anonymous):

h(d)=ad^2+bd+c (g is gravitational acceleration 63=-b/2a (the top point) b=-126a c=0 (at the initial point, the height is 0) 17=a(63)^2+ 63b 17=a*63^2-126*63a a=-17/(63^2) b= 34/63 h(d)=(-17/3969)d^2+34b/63

OpenStudy (aryavegapunk):

it would just be 34/63 and not 34b/63 right?

OpenStudy (aryavegapunk):

@marsss

OpenStudy (anonymous):

it's 34d/63. i'm so sorry.

OpenStudy (aryavegapunk):

It's alright, Thank you so much!!

OpenStudy (anonymous):

welcome my punker friend :)